### v1

# Training on the Test Task Confounds Evaluation and Emergence

Abstract
We study a fundamental problem in the evaluation of large language models that we call training on the test task. Unlike wrongful practices like training on the test data, leakage, or data contamination, training on the test task is not a malpractice. Rather, the term describes a growing set of techniques to include task-relevant data in the pretraining stage of a language model. We demonstrate that training on the test task confounds both relative model evaluations and claims about emergent capabilities. We argue that the seeming superiority of one model family over another may be explained by a different degree of training on the test task. To this end, we propose an effective method to adjust for training on the test task by fine-tuning each model under comparison on the same task-relevant data before evaluation. We then show that instances of emergent behavior largely vanish once we adjust for training on the test task. This also applies to reported instances of emergent behavior that cannot be explained by the choice of evaluation metric. Our work promotes a new perspective on the evaluation of large language models with broad implications for benchmarking and the study of emergent capabilities.222Code and fine-tuned models are available at https://github.com/socialfoundations/training-on-the-test-task
1 Introduction
The machine learning community has long recognized certain clear violations of the benchmarking protocol. Training on the test set is the most notorious among them (Duda and Hart, 1973; Hastie et al., 2017; Hardt and Recht, 2022). Data leakage (Kapoor and Narayanan, 2022) and data contamination (Roberts et al., 2023; Jiang et al., 2024) are closely related problems linked to the rise of massive web-crawled training datasets. Researchers can all agree that test data should never appear in the training set.
But it’s been much less clear what to do about legitimate attempts to bring training closer to evaluation. There is an obvious a gap between next token prediction at training time and tasks, such as reasoning and question answering, at test time. Ongoing research and engineering efforts, in fact, aim to narrow precisely this gap (MetaAI, 2024; Lewis, 2024). Why shouldn’t training be informed by knowledge about the downstream test tasks? What’s an unfair advantage of some may be the feature of others.
In this work, we group strategies to utilize task knowledge at training time under the umbrella term of training on the test task. Examples of training on the test task include the use of instruction-tuning data or question answering templates during pre-training (Zeng et al., 2022; Bai et al., 2023). We work from the premise that training on the test task is acceptable—or at least, unavoidable.
In a nutshell, we show that training on the test task strongly confounds model comparisons across different scales and model families. Moreover, it significantly obscures the study of emergent capabilities of large language models. Perhaps counterintuitively, we propose to mitigate the effects of training on the test task by doing more of it. We show that we can effectively level the playing field by giving each model the same, sufficient task-specific fine-tuning before evaluation. This adjustment restores clean log-linear scaling and makes capabilities predictable based on much smaller model scales.
1.1 Our contributions
We introduce the term training on the test task to group a growing repertoire of practices that utilize knowledge about evaluation tasks at training time. We study its impact on benchmark evaluations by inspecting 53 different language models in two major active benchmarks, MMLU and GSM8K.
We start in Section 2 by dividing models into those trained before November 2023 and those trained after. We find that for the same amount of compute, newer models outperform older models on average by 7 percentage points in MMLU and 17 points in GSM8K. We then fine-tune all models on the same amount of task-specific data before evaluation. We show that after fine-tuning, newer models no longer outperform older models. See Figure 1. This outcome suggests that newer models outperform older ones because they—implicitly or explicitly—trained more on the test task. Moreover, it shows how test task training can distort benchmark performance.
We propose a simple and effective method to adjust for the effect of training on the test task. Put simply, we fine-tune each model on the same, sufficient amount of task-specific data before evaluation. To validate our method, we demonstrate its effectiveness in a controlled setting: we take the older models and fine-tune half of them on the test task. Remarkably, this recreates the kind of performance differences observed between newer and older models. We then show that we can undo the advantage of the fine-tuned models over the other models by further fine-tuning all models on the test task (Section 3.1, Figure 3).
Next, we give evidence that training on the test task may be a more dominant factor in benchmark performance than data contamination. To argue this point, we consider ARC and HellaSwag, which use cloze prompts for evaluation. Here, at first there appears to be no sign of an unfair advantage in any specific model family. But after switching to MMLU-style multiple choice prompts, we see the same confounded results as for MMLU (Section 3.2, Figure 4). This suggest that newer models perform well in MMLU likely not because of memorization of specific testing data, but rather due to an improved ability to comprehend MMLU-style prompts. Either way, our proposed adjustment recovers fair model comparisons.
We show that training on the test task significantly distorts model family comparisons. The design choices of certain model families –such as its pretraining data mixture– may appear superior to others before adjusting for test task training but not after adjustment (Section 4.1, Figure 6). We also demonstrate that test task training overestimates the progress in capabilities achieved by recent models. After adjustment, newer models only modestly improve the Pareto frontier of performance against compute (Section 4.2, Figure 7).
Finally, we demonstrate that training on the test task has profound implications for the study of emergent capabilities. Specifically, we show that the phenomenon of emergence disappears gradually as the amount of training on the test task grows (Section 5). In particular, we can make capabilities visible and predictable from much smaller model scales. Importantly, our adjustment also works in cases, like MMLU, where previous purported explanations of emergence, such as the choice of evaluation metric, do not suffice.
Our work calls for a major reorientation of large language model evaluation. Model comparisons, scaling laws, and claims of emergence, are all strongly confounded by the choice of training data relative to the test tasks. Rather than scrambling to detect and disallow various forms of training on the test task, we propose to “fight fire with fire”. Specifically, our recommendation is to give each model the same sufficient amount of fine-tuning on task-relevant data prior to evaluation.
Limitations.
Although deliberate training on the test task can level the playing field, our work shows that generally significant fine-tuning is needed before the playing field is actually level. This requirement poses additional computational burden on the side of the evaluator. The computational resources to fine-tune all models equally may not be available to all evaluators depending on the circumstances. In addition, sufficient task-relevant training data might be expensive to source or generally unavailable for many tasks. While this is not an issue when model trainers also lack access to such data, training on proprietary task-specific training data would be difficult to correct for.
2 Adjusting for training on the test task
We choose MMLU (Hendrycks et al., 2020) and GSM8K (Cobbe et al., 2021) as a case study for investigating training on the test task in active benchmarks. MMLU tests for world knowledge, whereas GSM8K tests multi-step mathematical reasoning. These two benchmarks are very prominent in the literature at present time. For instance, GPT 4 (Achiam et al., 2023), Claude 3 (Anthropic, 2024), Gemini (Gemini et al., 2023) and Llama 3 (MetaAI, 2024) all report and highlight MMLU and GSM8K. They are also included in the HuggingFace (HF) Open LLM Leaderboard111See Appendix C for results pertaining to the OpenLLM Leaderboard v2 (Fourrier et al., 2024a). (Beeching et al., 2023), a popular benchmark leaderboard that evaluates and ranks models with publicly available weights. We evaluate models using LM Evaluation Harness library (EleutherAI, 2024), in identical fashion to the HF leaderboard.
We evaluate 53 base models, ranging in size from 70M to 70B parameters. See Appendix A.1 for the full list. The HF leaderboard’s FAQ makes the distinction between “base pretrained models” and instruction-tuned or chat models, arguing that this is necessary to ensure fair model comparisons. We select models that are categorized as “pretrained”. We check that the technical report of each of the selected models makes no mention of the model being fine-tuned. We only consider models for which the number of training tokens is known. This allows us to estimate the total amount of pretraining compute in FLOPs as , where is pretraining compute, is the number of model parameters, and is the number of training tokens.
Recent models outperform older ones given the same pretraining compute.
We evaluate models on MMLU and GSM8K, and plot benchmark accuracy against pretraining compute in Figure 1 top. We observe that performance correlates with pretraining compute for both benchmarks. However, on the surface it appears that more recent models better leverage pretraining compute. In other words, for a given compute budget newer models are able to attain better benchmark performance. In fact, models trained after November 2023 Pareto dominate those trained before November 2023.
These improvements in benchmark performance coincide with a recent trend in LLM research of increasingly utilizing test task knowledge at training time. For example, Qwen 1.5 (Bai et al., 2023), Olmo 1.7 (Groeneveld et al., 2024) and MAP Neo (Zhang et al., 2024) include instruction data during pretraining. StableLM 2 (StabilityAI, 2023) reformulates some of its pretraining datasets to better resemble downstream tasks such as question-answering. More subtly, the pretraining data mixture of Gemma (Gemma et al., 2024) was determined partially based on downstream benchmark evaluations.
This raises an important question: Do newer models outperform older ones mainly because newer models effectively trained more on the test task? At first sight, an answer seems elusive. After all, it would be both infeasible and cost prohibitive to train all models with the same training data and compute budget. Nevertheless, in the next section, we propose a way to get at the answer by adjusting for the effect of training on the test task.
2.1 Adjusting for training on the test task by training on the test task
We propose to adjust for differences in test task training by fine-tuning all models on the same, sufficient amount of task-specific data before evaluation. To do so, we need a source of task-specific data for each of the tasks we consider. For multiple choice questioning answering (MMLU), we use the auxiliary training set accompanying the HF MMLU repository222https://huggingface.co/datasets/cais/mmlu. It contains around 100,000 training examples and around 30M tokens. For mathematical reasoning (GSM8K), we combine the the MetaMathQA (Yu et al., 2023b) and Orca-Math (Mitra et al., 2024) datasets, totalling 600,000 training examples and approximately 200M tokens. We fine-tune models for three epochs using standard hyperparameter choices, with minimal hyperparameter tuning, see Appendix A.2. Note that the amount of compute required for fine-tuning is minimal in comparison to the compute required for pretraining, since all models considered were pretrained on at least 300B tokens.
We plot model scores on MMLU and GSM8K after fine-tuning in Figure 1 (bottom). We observe that after fine-tuning on task relevant data, newer models no longer Pareto dominate in terms of accuracy per pretraining compute. Instead, benchmark performance is strongly correlated with compute and both newer and older models follow remarkably similar scaling trends. That is, newer models no longer appear to outperform older models. Moreover, we observe that older models tend to benefit from training on the test task much more than newer models, see Figure 2. The improvements of older models are striking, often jumping from random chance accuracy to double digit improvements in accuracy. In contrast, fine-tuning brings comparatively little benefit to newer models. This observation suggests that newer models have already been trained on a substantial amount of task-relevant data.
2.2 Quantifying performance differences between newer and older models
We draw inspiration from scaling laws (Kaplan et al., 2020) in how we model benchmark accuracy to scale log-linearly with pretraining compute . To account for emergence (Wei et al., 2022), we assume that models perform at the task’s random chance accuracy up to scaling to some point of emergence . We let the variable denote whether a model was trained after November 2023, and regress the model
where , and are the fit’s parameters, and is random noise. We focus on the coefficient , which corresponds to the average difference in benchmark performance between newer and older models after controlling for pretraining compute. We fit the model in Equation 1, and report the regression coefficient in Figure 1. We obtain R for all model fits. Before adjusting for test task training, the estimated difference in performance between newer and older models are statistically significant, positive, and large. Specifically, recent models on average outperform older ones by over 7 accuracy points in MMLU and 17 accuracy points in GSM8K. These are remarkable differences in benchmark performance, as small single digit improvements are typically considered substantial improvements by the literature.
We repeat the analysis but using models’ adjusted benchmark scores, that is, those obtained after fine-tuning on the test task. After adjusting for test task training we find no evidence for a significant difference in benchmark performance between newer and older models. That is, the estimated coefficient is both small and not statistically significant. Put simply, newer models no longer outperform older ones. Therefore, conditioned on all models training on the same, sufficient amount of task-specific data before evaluation, there are no differences in benchmark performance between newer and older models.
Our findings provide evidence that the differences in benchmark performance between newer and older models are largely attributable to differences in test task training. We present a causal interpretation of our results in Appendix B, outlying the causal assumptions needed to establish that the improvements of newer models are attributable to training on the test task. Overall, we find no evidence for the improvements in performance of newer models being attributable to anything other than training more on the test task.
We include in Appendix B.1 a robustness check on the temporal split chosen, by instead diving models based on whether they were trained primarily on English language data. We obtain similar differences in performance, which we interpret as a valuable robustness check of our results. In Appendix C we instead consider the benchmarks of the newly released HF OpenLLM Leaderboard v2 (Fourrier et al., 2024a). Whereas the HF leaderboard v2 pays particular attention to guarding against data contamination (Fourrier et al., 2024b), we nonetheless find evidence that training on the test task confounds all benchmarks included in the Leaderboard v2. These findings highlight that training on the test task is a distinct phenomenon from data contamination, and new methods –such as our proposed adjustment procedure– are required to mitigate the confounding effect of training on the test task on benchmark evaluations.
3 Recreating differences in benchmark performance
Previously, we introduced a way to adjust for training on the test task. Here we systematically test the validity of this adjustment method. To do so, we demonstrate how to recreate the observed differences in performance between newer and older models by actively manipulating how much models train on the test task.
We do so in two ways. First, we fine-tune older models on task relevant data (Section 3.1). Second, we reformulate certain test tasks to use multiple choice prompts instead “cloze” evaluation (Section 3.2). Both experiments turn out to recreate the kind of performance difference we observed earlier. This not only provides further evidence that differences in performance between older and newer models are linked to test task training. It also demonstrates how test task training distorts benchmark evaluations.
Fortunately, in both cases, we show that fine-tuning models on task-relevant data before evaluation is an effective mechanism for mitigating the bias introduced by training on the test task.
3.1 Fine-tuning on the test task
For this section, we only consider models trained before November 2023, since we hypothesize that older models do not train on the test task much. We randomly split models into two cohorts: a control group and a treatment group. We fine-tune the treatment group on the datasets of task-relevant data introduced in Section 2. We fine-tune on each dataset independently, for a single epoch. We then evaluate the benchmark performance of the two cohorts, as well as their performance after adjusting for test task training. As in the previous section, we adjust for test task training by fine-tuning all models on the test task before evaluation.
We plot in Figure 3 the two cohorts’ benchmark performance before and after the adjustment. We repeat the statistical analysis of Section 2.2 and report the estimated coefficient indicating the average difference in benchmark performance between the two cohorts when controlling for compute.
Fine-tuning the treatment group results in large differences in performance between the control group and the treatment group, see Figure 3 middle. Qualitatively, the differences between the control and treatment group resemble those observed between newer and older models in Section 2.2. In particular, the fine-tuned models Pareto dominate the non fine-tuned models. Quantitatively, the estimated increase in performance due to fine-tuning is statistically significant and large. Importantly, it is also similar to the difference in performance between newer and older models estimated in Section 2.2. Therefore, fine-tuning older models on the test task gives rise to qualitatively and quantitatively similar confounding to that observed between newer and older models. This is consistent with our running hypothesis that newer models are largely equivalent to older models that trained on the test task.
After adjusting for test task training by further fine-tuning both the control and treatment groups on the test task, we observe that models in the treatment group are no longer outliers in terms of performance-per-compute, see Figure 3 right. Quantitatively, the estimated increase in performance is both small and not statistically significant. We therefore validate a vital soundness property of the proposed adjustment procedure: after deliberately training some models on the test task, we can undo their advantage over other models by further training all models on the test task.
3.2 Reformulating the test task
In this section we consider two additional benchmarks from the HF leaderboard: ARC Challenge (Clark et al., 2018) and HellaSwag (Zellers et al., 2019). Similarly to MMLU, ARC is comprised of grade-school level questions. HellaSwag instead tests for commonsense reasoning. Like MMLU, the questions in ARC and HellaSwag are accompanied by four possible answers, among which the model must differentiate the correct one. The standard MMLU evaluation formulates questions as multiple-choice: all four answer choices are listed, and the model is promoted to pick one. In contrast, ARC and HellaSwag use “cloze” evaluations: a models’ answer is taken to be that with the largest completion likelihood given the input question.
We evaluate all models on ARC and HellaSwag using the standard cloze evaluation, and plot their benchmark performance in Figure 4 left. We repeat the statistical analysis of Section 2.2, and report the average difference in performance between newer and older models after controlling for pretraining compute. Qualitatively, we observe that older models and newer models have very similar scaling trends. Quantitatively, the estimated difference in performance between newer and older models is small and not statistically significant. That is, newer models do not outperform older models on ARC and HellaSwag.
We then reformulate ARC and HellaSwag as MMLU-style multiple-choice questions, and plot the resulting benchmark performance in Figure 4 center. We observe large differences in performance between newer and older models. Qualitatively, these differences in performance resemble those observed for MMLU. In particular, newer models Pareto dominate in terms of performance-per-compute. Quantitatively, we find the difference in performance between newer and older models to be significant, positive, and large, and to be roughly similar in magnitude to that estimated for MMLU in Section 2.2. Therefore, reformulating the test task as multiple choice question answering leads to similar confounding to that observed for MMLU. This suggest that what causes the outliers in MMLU is likely not memorization of specific testing data (i.e., due to data contamination or leakage), but rather an improved ability for MMLU-style prompts.
We adjust for test task training by fine-tuning all models on the MMLU auxiliary training set, and plot their ARC Challenge and HellaSwag scores in Figure 4. We observe that newer models are no longer outliers in terms of performance-per-compute. Moreover, we no longer find evidence of a significant difference in performance between newer and older models. The proposed adjustment is therefore effective in removing the confounding resulting from newer models overperforming for MMLU-style prompts.
What does MMLU test for?
We evaluate MMLU using the “cloze” methodology instead of the usual multiple-choice prompts. We plot the results in Figure 5 center. With cloze evaluations, newer models are no longer are outliers in terms of MMLU performance. In fact, the difference in performance between newer and older models is now small and not statistically significant. This suggests that the standard MMLU evaluation conflates knowledge-testing with testing a models’ ability to answer multiple choice questions. For instance, smaller non fine-tuned models suffer from particularly strong A-bias for multiple-choice questions (Dominguez-Olmedo et al., 2023). Newer models therefore attain higher MMLU scores largely because they are better at multiple-choice question answering, and not because they necessarily “know more”.
4 Implications for model comparisons
Our findings indicate that training on the test task acts as a major confounder of LLM benchmark evaluations. We now discuss its implications for the relative comparison of model families (Section 4.1), as well as its implications for measuring progress in model capabilities over time (Section 4.2).
4.1 Comparing model families
We compare the MMLU and GSM8K performance of the Pythia, Llama 2, and Qwen 1.5 model families, which likely train on the test task to very different extents. Pythia was trained on the Pile (Gao et al., 2020), a collection of curated datasets that are unlikely to contain much test task data. Llama 2 was trained mostly on web data, which is reasonable to assume may contain test task data. Lastly, Qwen 1.5 explicitly includes instruction data in its pretraining mixture, thus likely training on the test task to a large extent.
In Figure 6 we plot the MMLU and GSM8K scores of the Llama 2, Qwen 1.5, and Pythia families of models, as well as their adjusted accuracy (i.e., after fine-tuning on task relevant data). Without adjustment, Qwen 1.5 appears to be the superior model family: it Pareto dominates both the Llama 2 and Pythia models. Furthermore, all Pythia models perform no better than random chance, and thus it is unclear what benefit scaling Pythia might bring. After fine-tuning the models on the the test task, however, all three model families exhibit very similar scaling trends. Therefore, when correcting for the confounding introduced by test task training it is unclear if any of the model families is superior to the others beyond their pretraining compute.
Interestingly, recent work equates pretraining data quality with downstream benchmark performance (Penedo et al., 2024; Li et al., 2024). For example, the Pile (Pythia’s pretraining dataset) is thought to be inferior to filtered web data. In light of our findings, it is plausible that a major contributing factor to the superior performance of “higher quality” pretraining datasets is that they contain a larger share of test task data.
4.2 Progress in model capabilities
One purpose of benchmarks is to track progress in model capabilities. In Figure 7 we plot the Pareto frontier of benchmark accuracy against pretraining compute, both for models trained before November 2023 and for all models. We measure progress by considering the area of improvement of the Pareto frontier since November 2023, shaded in green. Without adjustment, the difference between the two Pareto frontiers is rather large for both MMLU and GSM8K, indicating substantial progress since November 2023. After fine-tuning models on the test ask, however, the area of improvement reduces by a sixfold. Therefore, the confounding introduced by test task training leads to substantially overestimating the progress in MMLU and GSM8K capabilities per unit of compute achieved by recent model families.
On the other hand, recent models tend to be trained on more data. Given the Chinchilla scaling laws (Hoffmann et al., 2022), it is remarkable that newer, smaller models match the performance of older, larger ones for the same amount of pretraining compute. Since inference and fine-tuning of smaller models is substantially cheaper, recent models can be much more accessible to less well-resourced institutions, with little cost in performance. For example, we find that Llama 3 8B closely matches the performance of Llama 2 70B.
5 Implications for emergence
Throughout our evaluations, we observe emergent behaviour for MMLU and GSM8K: models perform at near random chance up to a certain scale of pretraining compute, followed by relatively sharper improvements in performance at larger scales (Wei et al., 2022). After training on the test task, however, emergence for MMLU and GSM8K appears to occur at substantially lower scales. We dedicate this section to more closely investigate the relationship between training on the test task and emergence.
Emergence arises at lower scales with increased test task training.
We consider only models trained before November 2023, as we have established that these models train on the test task to a lesser extent. We evaluate the models at intermediate checkpoints as they train on the datasets of task relevant data introduced in Section 2.1. We fit and in Equation 1 to the different intermediate checkpoints, and report in Figure 8 the corresponding points of emergence . We find that emergence arises at increasingly lower compute regimes as models increasingly train on the test task. For instance, for MMLU the non-fine tuned models exhibit emergence at around FLOPs, roughly the scale of Pythia 6.9B. After training on 64,000 examples, emergence arises around around FLOPs, that is, roughly the scale of Pythia 410M. That is, the benchmark performance of models after training on the test task is predictable at substantially lower scales.
Training on the test task yields increasingly better log-linear fits.
The log-linear relationship between pretraining loss and compute is well-established (Kaplan et al., 2020; Hoffmann et al., 2022). We observe that, for the compute ranges that we consider, training on the test task increasingly recovers log-linear scaling between pretraining compute and benchmark accuracy. Similarly to the earlier section, we evaluate intermediate checkpoints but instead fit log-linear functions in Figure 9. We observe that the R2 of the fit improves substantially as the models train on more task-relevant data. For MMLU, the value jumps from to after training on 64,000 examples. Therefore, after training on the test task almost all of the variation in accuracy can be explained by log-linear scaling of pre-training compute.
Discussion.
Schaeffer et al. (2024a) argue that emergence appears due to the choice of metric. To mitigate emergence, they suggest to consider Brier score instead of accuracy. We observe, however, that the emergent behaviour of MMLU does not disappear when using the Brier score, see Figure 5 right, nor that of ARC and HellaSwag when framed as multiple-choice questions, see Figure 16 in Appendix D. While more complex changes of metric might resolve the emergence in multiple-choice QA (Schaeffer et al., 2024b), we discuss two practical solutions to obtain predictive scaling while maintaining accuracy as the evaluation metric.
For MMLU and multiple-choice benchmarks more broadly, we consistently observe that cloze evaluations yield smoother and more predictable scaling even when using accuracy as the evaluation metric. Since the purpose of these benchmarks is knowledge-testing more so than testing multiple-choice answering ability, cloze evaluations should be preferable insofar predictive scaling is an important consideration.
More broadly, if sufficient task relevant data is available, then training on the test task can result in much more predictable scaling by shifting emergence to smaller compute scales. Crucially, the evaluation metric and methodology need not be changed. Note that in many settings it is not a priori apparent what metric or evaluation methodology might results in predictive scaling. Scaling laws after fine-tuning correspond to those of more “specialist” models, which for some domains –such as the legal domain (Dominguez-Olmedo et al., 2024)– or purposes –e.g., safety– might be preferable to the scaling law of generalist models.
6 Related work
Benchmarks have played a central role in both machine learning (Hardt and Recht, 2022) and natural language processing (Storks et al., 2019). Classically, benchmarks comprised both a test set and a reasonably large training set (Garofolo et al., 1993; LeCun et al., 1998; Sang and De Meulder, 2003; Koehn, 2005; Deng et al., 2009). Models were trained on the same training set, and then evaluated on the accompanying test set. The success of unsupervised language modelling (Peters et al., 2018; Kenton and Toutanova, 2019; Radford et al., 2019), however, has changed this paradigm. Firstly, present-day language models differ in their training data, which is not standardized but rather treated as a design choice (Raffel et al., 2020; Albalak et al., 2024; Li et al., 2024). Secondly, language models are a priori not trained with the explicit objective of maximizing any single benchmark score. Rather, language models are expected to be able to perform a broad range of tasks (Wang et al., 2018; Brown et al., 2020). Consequently, models are evaluated and compared using a plurality of benchmarks (Beeching et al., 2023; Liang et al., 2023; Srivastava et al., 2023).
Data contamination.
Data contamination or test-set contamination refers to any overlap between the training and the test data such that test results overestimate a model’s generalization performance. The scale and often little curation of present-day pretraining corpora exacerbates data contamination concerns in language model evaluations (Jiang et al., 2024). Consequently, data contamination is usually discussed in the technical reports accompanying model releases (Radford et al., 2019; Brown et al., 2020; Chowdhery et al., 2023; Touvron et al., 2023b). However, detecting and preventing data contamination is currently an open problem (Gunasekar et al., 2023; Yang et al., 2023b; Golchin and Surdeanu, 2023). Roberts et al. (2023) and Li and Flanigan (2024) find that models often perform better on datasets that were publicly available during model training. While almost all models that we consider were released after MMLU and GSM8K, we nonetheless find that, controlling for compute, more recent models perform better. These performance gains are unlikely to be driven solely by test set leakage and require additional explanation.
Training on the test task.
The effectiveness of fine-tuning on the training set accompanying LLM benchmarks is well-known (Wei et al., 2021; Wang et al., 2022; Chung et al., 2024). Consequently, many influential instruction-tuning datasets contain or are partly derived from benchmark train data (Wei et al., 2021; Honovich et al., 2022; Mukherjee et al., 2023). Li and Flanigan (2024) identify small amounts of benchmark-specific data in the publicly available Alpaca (Taori et al., 2023) and Vicuna (Chiang et al., 2023) instruction-tuning sets. Zhou et al. (2023b) empirically analyze the effects of fine-tuning on benchmark-specific data and warn about its impacts on benchmark validity. To circumvent these issues, recent work has focused on indirect indicators of broader data contamination, such as a lack of robustness to task transformations (Wu et al., 2023), or underperformance on benchmarks with novel task combinations (Yu et al., 2023a). In contrast, we find evidence for training on the test task without the need for explicitly identifying specific data points used at training time, or modifying tasks. In addition, our method allows us to quantify and correct for the effects of training on the test task on benchmark performance.
Emergent abilities of language models.
Emergent capabilities (Wei et al., 2022; Ganguli et al., 2022) refer to levels of model performance at large scales that cannot be easily predicted by extrapolating from smaller scales. Wei et al. (2022) report emergent capabilities for various benchmarks including MMLU and GSM8K (Srivastava et al., 2022). However, Srivastava et al. (2022); Schaeffer et al. (2024b) find that the log-probability of the correct answer often improves smoothly, even when other metrics seem to show emergence. Lu et al. (2023) argue that most emergent capabilities can be explained by in-context-learning. Schaeffer et al. (2024a) argue that emergent capabilities are mostly an artifact of non-linear and discontinuous evaluation metrics like accuracy. In contrast, we find signs of emergence on tasks like MMLU, even when using continuous metrics like the Brier score. We additionally show that fine-tuning on the test task yields more predictive scaling by shifting the point of emergence to substantially smaller compute scales.
7 Discussion
The 1968 Olympics took place in Mexico City at the significant altitude of 2340 meters, higher than Australia’s tallest peak. Runners who had trained at altitude in their home countries were better prepared to compete in Mexico City’s conditions, as it turned out. But the hotly debated results of the Games did not lead the organizers to prohibit training at natural altitude. Instead, they let everyone do it; and athletes came to consider altitude training an excellent way to train.
The anecdote holds a lesson for the evaluation of large language models half a century later. Knowledge about the evaluation conditions necessarily influences training practices under competitive pressure. It may be a fool’s errand to prohibit the practice. Instead, we propose to adjust for it by giving every model the same task-specific preparation before evaluation. We work from the assumption that training on the test task, in general, cannot be effectively detected, disallowed, or disincentivized. Detecting what training data a model has seen is a notoriously difficult problem—existing heuristics achieve partial success at best. Researchers routinely acknowledge the futility of fighting data contamination. Moreover, we anticipate that the ways to effectively train on the test task will only grow in scope and adoption.
Our work demonstrates that comparisons of different models are confounded by the choice of training data and training practices. Different model families vary in the degree that they were—implicitly or explicitly—trained on various test tasks. It therefore makes little sense to compare model performance at face value without accounting for how the training data relate to the test task. The same problem extends to scaling. Smaller models can appear unexpectedly performant if they were trained to a greater extent on task data.
We can apply the same principles to emergent behavior. After training on the test task, model capabilities become predictable at smaller model size and grow continuously with scale. This is not to say that emergence isn’t real; it may well be a real phenomenon for a fixed choice of dataset and evaluation metric. But training on the test task removes the unpredictability and discontinuity associated with emergence, notably without any change in the metric, thus largely disarming the ominous nature of emergence.
Despite the daunting challenges that training on the test task poses for the fair evaluation of language models, it’s also its own best remedy. Giving each model the same sufficient task-specific fine-tuning harmonizes model comparisons, deconfounds scaling laws, and linearizes the relationship between model capabilities and log-scale. We hope that our work informs stronger evaluation standards that address central challenges in the current evaluation ecosystem. Our proposal has the added side benefit of creating incentives for model builders to create models that can be fine-tuned easily and respond well to fine-tuning.
Appendix A Additional experimental details
A.1 Models considered
Model size in billions of parameters is indicated by and pretraining data size in trillions of tokens is indicated by . Model weights were retrieved from the corresponding HuggingFace (HF) repositories.
| Name | Train date | HF repository | Citation | ||
| baichuan-13b | 2023-06 | 13 | 1.4 | baichuan-inc/Baichuan-13B-Base | Yang et al. (2023a) |
| baichuan-7b | 2023-06 | 7 | 1.2 | baichuan-inc/Baichuan2-7B-Base | Yang et al. (2023a) |
| baichuan2-13b | 2023-09 | 13 | 2.6 | baichuan-inc/Baichuan2-13B-Base | Yang et al. (2023a) |
| baichuan2-7b | 2023-09 | 7 | 2.6 | baichuan-inc/Baichuan2-7B-Base | Yang et al. (2023a) |
| falcon-11b | 2024-05 | 11 | 5.0 | tiiuae/falcon-11B | Almazrouei et al. (2023) |
| falcon-7b | 2023-04 | 7 | 1.5 | tiiuae/falcon-7b | Almazrouei et al. (2023) |
| gemma-2b | 2024-02 | 2 | 2.0 | google/gemma-2b | Gemma et al. (2024) |
| gemma-7b | 2024-02 | 7 | 6.0 | google/gemma-7b | Gemma et al. (2024) |
| gpt-j-6b | 2021-03 | 6 | 0.4 | EleutherAI/gpt-j-6b | Wang and Komatsuzaki (2021) |
| internlm-20b | 2023-09 | 20 | 2.3 | internlm/internlm-20b | InternLM (2023) |
| internlm-7b | 2023-07 | 7 | 1.0 | internlm/internlm-7b | InternLM (2023) |
| internlm2-base-20b | 2024-01 | 20 | 2.6 | internlm/internlm2-base-20b | Cai et al. (2024) |
| internlm2-base-7b | 2024-01 | 7 | 2.6 | internlm/internlm2-base-7b | Cai et al. (2024) |
| llama-13b | 2023-02 | 13 | 1.0 | None | Touvron et al. (2023a) |
| llama-2-13b | 2023-07 | 13 | 2.0 | meta-llama/Llama-2-13b-hf | Touvron et al. (2023b) |
| llama-2-70b | 2023-07 | 70 | 2.0 | meta-llama/Llama-2-70b-hf | Touvron et al. (2023b) |
| llama-2-7b | 2023-07 | 7 | 2.0 | meta-llama/Llama-2-7b-hf | Touvron et al. (2023b) |
| llama-3-8b | 2024-04 | 8 | 15.0 | meta-llama/Meta-Llama-3-8B | MetaAI (2024) |
| llama-30b | 2023-02 | 32.5 | 1.4 | None | Touvron et al. (2023a) |
| llama-65b | 2023-02 | 65.2 | 1.4 | None | Touvron et al. (2023a) |
| llama-7b | 2023-02 | 7 | 1.0 | None | Touvron et al. (2023a) |
| map-neo-7b | 2024-05 | 7 | 4.5 | m-a-p/neo_7b | Zhang et al. (2024) |
| olmo-1.7-7b | 2024-04 | 7 | 2.05 | allenai/OLMo-1.7-7B-hf | Groeneveld et al. (2024) |
| olmo-1b | 2024-01 | 1 | 2.0 | allenai/OLMo-1B-hf | Groeneveld et al. (2024) |
| olmo-7b | 2024-01 | 7 | 2.46 | allenai/OLMo-7B-hf | Groeneveld et al. (2024) |
| openllama-13b | 2023-06 | 13 | 1.0 | openlm-research/open_llama_13b | OpenLlama (2023) |
| openllama-3b | 2023-06 | 3 | 1.0 | openlm-research/open_llama_3b | OpenLlama (2023) |
| openllama-3b-v2 | 2023-07 | 3 | 1.0 | openlm-research/open_llama_3b_v2 | OpenLlama (2023) |
| openllama-7b | 2023-06 | 7 | 1.0 | openlm-research/open_llama_7b | OpenLlama (2023) |
| openllama-7b-v2 | 2023-07 | 7 | 1.0 | openlm-research/open_llama_7b_v2 | OpenLlama (2023) |
| pythia-1.4b | 2023-02 | 1.4 | 0.3 | EleutherAI/pythia-1.4b | Biderman et al. (2023) |
| pythia-12b | 2023-02 | 12 | 0.3 | EleutherAI/pythia-12b | Biderman et al. (2023) |
| pythia-160m | 2023-02 | 0.16 | 0.3 | EleutherAI/pythia-160m | Biderman et al. (2023) |
| pythia-1b | 2023-02 | 1 | 0.3 | EleutherAI/pythia-1b | Biderman et al. (2023) |
| pythia-2.8b | 2023-02 | 2.8 | 0.3 | EleutherAI/pythia-2.8b | Biderman et al. (2023) |
| pythia-410m | 2023-02 | 0.41 | 0.3 | EleutherAI/pythia-410m | Biderman et al. (2023) |
| pythia-6.9b | 2023-02 | 6.9 | 0.3 | EleutherAI/pythia-6.9b | Biderman et al. (2023) |
| pythia-70m | 2023-02 | 0.07 | 0.3 | EleutherAI/pythia-70m | Biderman et al. (2023) |
| qwen-1.5-0.5b | 2024-01 | 0.5 | 2.4 | Qwen/Qwen1.5-0.5B | Bai et al. (2023) |
| qwen-1.5-1.8b | 2024-01 | 1.8 | 2.4 | Qwen/Qwen1.5-1.8B | Bai et al. (2023) |
| qwen-1.5-14b | 2024-01 | 14 | 4.0 | Qwen/Qwen1.5-14B | Bai et al. (2023) |
| qwen-1.5-4b | 2024-01 | 4 | 2.4 | Qwen/Qwen1.5-4B | Bai et al. (2023) |
| qwen-1.5-7b | 2024-01 | 7 | 4.0 | Qwen/Qwen1.5-7B | Bai et al. (2023) |
| redpajama-3b | 2023-05 | 3 | 0.8 | togethercomputer/RedPajama-INCITE-Base-3B-v1 | TogetherWeCompute (2023) |
| redpajama-7b | 2023-05 | 7 | 1.0 | togethercomputer/RedPajama-INCITE-7B-Base | TogetherWeCompute (2023) |
| skywork-13b | 2023-10 | 13 | 3.2 | Skywork/Skywork-13B-base | Wei et al. (2023) |
| stablelm-2-1.6b | 2024-01 | 1.6 | 2.0 | stabilityai/stablelm-2-1_6b | Bellagente et al. (2024) |
| stablelm-2-12b | 2024-03 | 12.1 | 2.0 | stabilityai/stablelm-2-12b | Bellagente et al. (2024) |
| stablelm-3b-4e1t | 2023-09 | 2.8 | 4.0 | stabilityai/stablelm-3b-4e1t | StabilityAI (2023) |
| stablelm-base-alpha-3b-v2 | 2023-08 | 2.8 | 1.1 | stabilityai/stablelm-base-alpha-3b-v2 | StabilityAI (2023) |
| stablelm-base-alpha-7b-v2 | 2023-08 | 7 | 1.1 | stabilityai/stablelm-base-alpha-7b-v2 | StabilityAI (2023) |
| yi-6b | 2023-11 | 6 | 3.0 | 01-ai/Yi-1.5-6B | Young et al. (2024) |
| ziya2-13b-base | 2023-11 | 13 | 2.65 | IDEA-CCNL/Ziya2-13B-Base | Gan et al. (2023) |
A.2 Fine-tuning hyperparameters
We fine-tune all model parameters. For models with less than B parameters, we fine-tune on a single GPU with BF16 precision. For models between B and B parameters, we train on a single H100 node using DeepSpeed ZeRO-3 (Rajbhandari et al., 2020) and full precision. For models with more than B parameters, we train on two H100 nodes using DeepSpeed ZeRO-3 and full precision. Due to the large compute cost of the experiments, we perform minimal hyperparameter tuning and use standard hyperparameter choices throughout. We use a learning rate of for models with fewer than B parameters and a learning rate of for models with more than B parameters. For four of the B models –Gemma 7B, Olmo 7B, Olmo 1.7 7B, and Llama 3 8B– benchmark accuracy heavily degraded after fine-tuning. For these models, we use a peak learning rate of instead. These four models were all released after November 2023. We use a cosine learning rate schedule with linear warm-up for 50 steps and decay to of the peak learning rate. We use AdamW (Loshchilov and Hutter, 2018) as the optimizer, with , , and . We fine-tune with batch size 64. We use a weight decay rate of 0.1 and clip gradients at 1.0. We verify that the training loss decreases for all models on both of the fine-tuning datasets. To reduce the computation burden of fine-tuning, we train with context size 600. We verify that less than 5% of the fine-tuning examples have context length above 600.
We use an internal cluster of A100 and H100 GPUs. Fine-tuning all models required approximately 10,000 H100 GPU hours, whereas evaluating all models in the different benchmarks required approximately 400 H100 GPU hours.
Appendix B Causal interpretation of our findings
In Section 2.2 we demonstrated that models trained after November 2023 significantly outperform those trained before November 2023 for both MMLU and GSM8K. We now seek to determine how much of the benchmark improvements of newer models is attributable to newer models training more on the test task. That is, the extent to which the effect of model recency on benchmark accuracy is mediated by training on the test task . The key obstacle to our analysis is that test task training is unobservable. Firstly, because practitioners are typically not transparent about their designs choices (e.g., pretraining data). Secondly, because the extent to which different training practices might amount to test task training is unclear. However, we are able to intervene on by fine-tuning on the test task.
Figure 10 summarizes our causal assumption. The time at which a model was trained determines the design choices made, such as its pretraining data or pretraining compute . These design choices in turn affect how much the model trains on the test task. All these factors ultimately influence the pretrained model and thus its benchmark performance. We assume that test task training does not causally influence pretraining compute, but compute might influence test task training. For instance, training on larger datasets may lead to training more on the test task.
We intervene on test task training by fine-tuning all models on the same, sufficient amount of task-specific data before evaluation. That is, we apply the adjustment proposed in Section 2.1. The external validity of our analysis hinges on the assumption that our experimental setting –fine-tuning models after the pretraining stage– is reasonably similar to the natural settings in which practitioners might train on the test task during pretraining (e.g., by including instruction data in the pretraining data mixture). We provide evidence in Appendix B.2 that this is the case.
We model fine-tuning as a hard intervention (Pearl, 2009). The specific magnitude of the intervention need not be quantified. Instead, the key assumption is that by fine-tuning on the same, sufficient amount of task data, all models will have received the same amount of test task training. Since some base models may have already trained on the test task prior to fine-tuning, our assumption will only hold if test task training saturates and we train on enough task data to reach saturation. We find evidence of saturation for both MMLU and GSM8K, see Appendix B.3.
We draw inspiration from scaling laws (Kaplan et al., 2020) and model relationship between pretraining compute and its causal descendants as pice-wise log-linear:
For simplicity, we consider three fixed knots at , , and FLOPs. We assume all other variable relationships to be linear, resulting in the structural assignments:
We denote benchmark accuracy after fine-tuning as . To estimate the direct effect of model recency on accuracy, we regress the linear model
where are the fit’s parameters and is random noise. The coefficient corresponds to the direct effect of model recency on benchmark accuracy. We additionally regress on the difference in accuracy pre and post intervention
where , , are the fit’s parameters and is random noise. The coefficient corresponds to the indirect effect of model recency on benchmark accuracy mediated by test task training (Pearl, 2013). That is, the improvements in accuracy of recent models attributable to training on the test task.
| MMLU | GSM8K | |
| -0.004 | 0.000 | |
| (0.009) | (0.032) | |
| R2 | 0.926 | 0.763 |
Standard errors in parentheses. Bold indicates .
| MMLU | GSM8K | |
| 0.071 | 0.168 | |
| (0.018) | (0.032) | |
| R2 | 0.530 | 0.503 |
Standard errors in parentheses. Bold indicates .
We fit the models in Equation 5 and Equation 6, and we report the coefficients pertaining to and in Table 3 and Table 3, respectively. We find no evidence of a significant direct effect of model recency on accuracy. On the other hand, its indirect effect mediated by test task training is significant, positive, and large.
Therefore, our analysis indicates that the differences in MMLU and GSM8K performance between newer and older models observed in Section 2.1 are primarily attributable to differences in test task training. That is, the mechanism by which newer models outperform older models is by training more on the test task.
B.1 Robustness check on the temporal split: EN vs CN language data
Instead of diving models using a temporal split, we divide models based on whether they were trained primarily on English (EN) data or on a mixture of English and Chinese (EN+CN) language data. While there is a considerable overlap between the temporal split and the EN/EN+CN model split, there are notable differences. In particular, the Baichuan, Baichuan 2, and InternLM, and Skywork families were trained before November 2023 and trained on EN+CN data. Conversely, Gemma, Llama 3, StableLM 2, Falcon 2, and Olmo were trained after November 2023 and trained on EN data.
We repeat the analysis of Section 2 for the EN and EN+CN model split. We observe that, controlling for pretraining compute, models trained on EN+CN language data outperform those trained primarily on EN by 9 accuracy points on MMLU and 12 accuracy points on GSM8K. After the proposed adjustment, however, the difference in performance between models trained on EN data and EN+CN data is small and not statistically significant.
The confounding and measured effect sizes for the EN and EN+CN model split resemble those obtained for the temporal split, which we interpret as a valuable robustness check of our results.
B.2 How similar are newer models to older, fine-tuned models?
In Section 3.1 we fine-tune older models on the test task and we demonstrate that the differences in benchmark performance between the fine-tuned and non fine-tuned models resemble those between newer and older models. In this section we provide further evidence that newer models resemble older, fine-tuned models.
We take the older models and we fine-tune them with 64,000 training examples from the auxiliary training sets introduced in Section 2.1. We plot in Figure 12 the benchmark scores of the older, fine-tuned models as well as that of the newer models. We qualitatively observe that both the older, fine-tuned models and the newer models exhibit similar scaling. That is, older fine-tuned models resemble newer models in terms of performance per compute.
We perform a quantitative analysis consisting in discriminating between the older models and the newer models based on their pretraining compute and benchmark accuracy. That is, we construct a tabular dataset where rows are models and columns are their corresponding pretraining compute, benchmark accuracy, and whether the model was trained after November 2023. We then train a classifier aiming to predict model recency from compute and accuracy. Intuitively, if the performance of older models is very different form that of newer models, then we would obtain high prediction accuracy (i.e., the two classes are highly separable). Note that prediction accuracy provides a lower bound on the total variation (TV) distance between the distributions of compute and accuracy of older and newer models.
| Discriminator test | MMLU | GSM8K |
| Older models vs | 64.6% | 73.9% |
| newer models | ||
| Fine-tuned, older models vs | 52.2% | 52.5% |
| newer models |
Random chance accuracy is 50%.
We train XGBoost classifiers and report balanced accuracy for leave-one-out cross-validation in Table 4. We find that newer models are reasonable distinguishable from older models, with 63% accuracy for MMLU and 79% accuracy for GSM8K. In contrast, we obtain close to random-chance accuracy in discriminating between older, fine-tuned models and newer models. That is, older fine-tuned models are indistinguishable from newer models in terms of their performance.
B.3 Test task training and saturation
We show that training on the test task saturates. We consider the intermediate checkpoints of the adjustment procedure, that is, fine-tuning for 3 epochs of the task datasets introduced in Section 2.1. We plot in Figure 13 the gain in benchmark accuracy from the first 75% (MMLU) and 80% (GSM8K) training steps, as well as the gain in benchmark accuracy from the remaining steps. Almost all of the performance improvements occur in the earlier training steps. In contrast, the final 25% (MMLU) and 20% (GSM8K) optimization steps result in almost no changes in benchmark accuracy. This indicates that training on the test task saturates, and we train for enough steps to reach saturation.
Note that while more training data might result in further benchmark improvements, we show that the task datasets that we use are sufficient for older models to reach the performance of newer models, see Appendix B.2.
Appendix C Results for the OpenLLM Leaderboard v2
HuggingFace released on June 2024 a revision of the OpenLLM Leaderboard (Fourrier et al., 2024a). The HF leaderboard v2 differs from v1 in the six benchmarks it considers: MMLU Pro (Wang et al., 2024), GPQA (Rein et al., 2023), BBH (Suzgun et al., 2023), MuSR (Sprague et al., 2023), the Level 5 subset of MATH (Hendrycks et al., 2021), and IFEval (Zhou et al., 2023a). MMLU and GPQA test for knowledge and are framed as multiple-choice questions. BBH and MuSR test for reasoning. MATH tests for mathematical reasoning. IFEval tests the ability of models to follow instructions.
The creators of the OpenLLM Leaderboard cite contamination as a key motivation for releasing the v2 revision. They note that a key criteria in choosing the benchmarks of the HF leaderboard v2 was lack of contamination in models as of today. In particular, Fourrier et al. (2024b) claim that current models are not contaminated for GPQA, MuSR, and MMLU Pro: GPQA due to the gating of the test set, and MuSR and MMLU Pro due to their “youth”. Fourrier et al. (2024b) succinctly express their concern as regards to data contamination in the HF leaderboard v1:
"Some newer models also showed signs of contamination. By this, we mean that models were possibly trained on benchmark data or on data very similar to benchmark data. As such, some scores stopped reflecting the general performance of the model and started to overfit on some evaluation datasets instead of reflecting the more general performance of the task being tested. This was, in particular, the case for GSM8K and TruthfulQA, which were included in some instruction fine-tuning sets."
Note that “models were possibly trained on benchmark data or on data very similar to benchmark data” encompasses not only test set contamination but more broadly training on the test task.
Bold indicates statistical significance with p < 0.05.
We evaluate all 53 models on MMLU Pro, GPQA, BBH, MuSR and MATH Lvl 5. We use the LM Evaluation Harness library in identical fashion to the HF leaderboard v2. We do not evaluated on IFEval since it tests for instruction following and we evaluate base models. We additionally evaluate the models that we fine-tuned in Section 2.1 for multiple choice question answering and mathematical reasoning. This gives us models’ adjusted benchmark scores after training on multiple choice question answering and mathematical reasoning. For MATH Lvl 5, we use the models fine-tuned on mathematical data, whereas for MMLU Pro, GPQA, BBH and MuSR we use the models fine-tuned on multiple choice question answering. The fine-tuning datasets were not adapted to the new benchmarks in the HF leaderboard v2, thus giving a valuable insight into how well these task-relevant datasets generalize beyond MMLU and GSM8K.
We plot in Figure 14 models benchmark scores pre and post post adjustment. We find that newer models significantly outperform older ones in all five benchmarks after controlling for pretraining compute. The differences in performance are smaller in absolute terms than those measured for MMLU (0.068) and GSM8K (0.168). This is in part because these benchmarks are “harder”, meaning also smaller differences in performance between the best and worst model. For this reason, we also report the difference between newer and older models relative to the difference between the best and worst model. This relative difference is 13.7% for MMLU Pro, 14.5% for GPQA, 12.1% for MuSR, 9.7% for BBH, and 10.0% for MATH Lvl 5, compared to 15.3% for MMLU and 25.0% for GSM8K. Therefore, newer models overperform in MMLU Pro, GPQA and MuSR about as much as they do for MMLU, and somewhat less for BBH and MATH Lvl 5.
Fine-tuning on task-relevant data reduces the difference in performance between newer and older models for all five benchmarks. For GPQA and MuSR, the difference in performance after adjustment is small () and not statistically significant. For BBH, the estimated difference in performance reduces by 40% to 0.015 and is no longer statistically significant. For MMLU Pro and MATH Lvl 5 the difference reduces by 19% and 33% respectively but remains reasonably large ( > 0.01) and statistically significant. Therefore, find evidence that training on the test task plays a substantial role in newer models outperforming older ones in the benchmarks of the HF Leaderboard v2.
Bold indicates statistical significance with .
One possible reason for the fact that the adjustment for MMLU Pro and MATH Lvl 5 is not as effective as for MMLU and GSM8K is that the fine-tuning examples are simply not as relevant for MMLU Pro and MATH Lvl 5. For example, the questions in MATH Lvl 5 contain much more LaTeX equation formatting than our mathematical reasoning fine-tuning dataset. Note that the answers to many MATH Lvl 5 questions are precisely formatted as LaTeX equations. Regarding MMLU Pro, our multiple choice fine-tuning dataset contains mostly questions with 4 answer choices, whereas all MMLU Pro questions have 10 answer choices. Thus, models are primarily fine-tuned to answer “A”, “B”, “C”, and “D” but not “E”, “F”, “G”. We modify MMLU Pro to only contain questions with 4 answer choices by for every question randomly discarding 6 of the incorrect answer choices. We evaluate models pre and post adjustment and plot the results in Figure 15. We observe that the difference in performance between newer and older models after adjustment reduces from 0.024 to 0.016, and is no longer statistically significant. This observation suggests that fine-tuning one more relevant task-data might further reduce the gap between newer and older models.
Discussion.
Fourrier et al. (2024b) cite newer models overperforming in the HF leaderboard v1 due to being “possibly trained on benchmark data or on data very similar to benchmark data” as a major reason for the HF leaderboard v2 revision. We however find evidence that training on the test task is also a confounder for the newly included benchmarks. Specifically, the difference in performance between newer and older models is significant for MMLU Pro, GPQA, MuSR, BBH and MATH Lvl 5, and these differences reduce after adjusting by fine-tuning on the test task.
Fourrier et al. (2024b) explicitly highlight GPQA and MuSR as benchmarks likely unaffected by contamination, the former due to being gated and latter due to its “youth”. Not only do newer models significantly outperform older ones in GPQA and MuSR, but these differences in performance fully vanish after fine-tuning on the test task. That is, newer models likely overperform in GPQA and MuSR precisely due to training on the test task.
These findings highlight that training on the test task is a distinct phenomenon from test set leakage. Strategies that aim to mitigate data contamination –e.g., dynamic benchmarks– might not be effective in mitigating the confounding effect of training on the test task. In contrast, we extensively demonstrated the effectiveness of our proposed adjustment procedure, that is, fine-tuning on sufficient task-relevant data before evaluation.
Appendix D Additional figures
In Figure 16 we show that ARC and HellaSwag do not exhibit emergence when using the standard cloze evaluation. When reformulating the task as multiple choice in the style of MMLU, however, we observe emergence around to FLOPs, similarly to MMLU. Emergence in this range of compute persists even when changing the evaluation metric from accuracy to Brier score –a continuous metric–, as suggested by Schaeffer et al. (2024a).